OpenStudy (anonymous):

find the corresponding quadratic equation of these pair of roots: {5/3, 1/2}

OpenStudy (anonymous):

since the roots are given, multiply out (x-5/3)(x-1/2)

OpenStudy (anonymous):

do u FOIL

OpenStudy (anonymous):

let roots be m and n so m=5/3 and n=1/2 let the quadratic equation be ax^2+bx+c=0 so m+n=sum of roots =-b/a=5/3+1/2=13/6 mn=product of roots =c/a=5/6 so